if AF = 2(x + 1)and AC = 5(x -2), find FC
Answer by checkley77(12569)

AF=2(x+1)
F=2(X+1)/A
AC=5(x-2)
C=5(X-2)/A
FC=[2(X+1)/A][5(X-2)/A]
FC=[(2X+2)/A][(5X-10)/A]
FC=10X^2+10X-20X-20)/A^2
FC=(10X^2-10X-20)/A^2
FC=10(X^2-X-2)/A^2
FC=10(X-2)(X+1)/A^2